# When Do Transformers Learn Heuristics for Graph Connectivity?

### v1

@@ 4.2 Expressivity and Capacity
If a 2-layer transformer fails to generalize in §3.3, should we attribute this to the architecture’s expressivity? We argue not. Theorem 4.3 shows that an $L$-layer disentangled transformer can implement the matrix powering algorithm and is perfect on graphs of diameter at most $3^{L}$. Moreover, Theorem 4.4 shows this $3^{L}$ threshold is tight and exact. To make this precise, we first formalize graph distance and diameter in Section 4.2.
@@
Throughout, we define the diameter of a graph, denoted $\mathrm{diam}(G)$, to be the maximum diameter among its connected components. Note this differs from the common convention on disconnected graphs, where the latter sets $\max_{u,v}d_{G}(u,v)=\infty$. Ours is always finite.
@@
Theorem 4.3 (Expressivity).
@@
For all $\ell$, setting $W_{\ell}=I_{d_{\ell-1}}$ suffices. These choices of weights implements the matrix powering algorithm $\sum_{j=0}^{\operatorname{diam}(G)}\alpha_{j}A^{j}$ with nonnegative coefficients $\alpha_{j}>0$. ∎
The expressivity result shows what is possible, we next show a capacity bound which reveals the model’s inherent limitations. We now prove a tight, non-asymptotic upper bound on the graph diameter an $L$-layer model can handle, linking model depth directly to instance difficulty.
Theorem 4.4 (Capacity).
@@
Case 1 (False positive occurs Section B.2). Suppose for some graph $H$ and a layer $\ell$, a positive score appears at $(u,v)$ in different components. Take $\ell$ minimal (the earliest layer across all graphs where a false positive emerges). We will create a new graph $G$ that (i) preserves this false positive on $(u,v)$ and (ii) contains a path of length $3^{L}$. To do so, we backtrack the computation DAG, tracing the “sources” that contribute to the false positivity of $(u,v)$. This gives us two subgraphs (one for $u$, one for $v$) with roots at layer $\ell-1$ that we call certificates. Because $\ell$ is minimal, every previous layer is free of false positives; hence, the two certificates induce two disjoint sets of graph nodes. Finally, create a new graph $G$ which embeds these two certificates and arranges all other nodes into a long chain. Then $G$ has the properties we seek.
Case 2 (No false positives Section B.2). Suppose now that no intermediate layer has false positives. We show that “information” spreads no faster than $3^{L}$ so that it never predicts “Yes” on node pairs with distance beyond $3^{L}$. We first apply the no-false-positives assumption on the empty (self-loops-only) graph. Inductively, each column of each hidden states is supported on exactly one row, which ranges from $1$ to $n$. This naturally gives a “label” for each column in each hidden states. The crux of the proof is to inductively show that at layer $\ell$, two columns can “share” information, thereby creating a positive score on $(u,v)$, only if their labels, interpreted as graph nodes, are within distance $3^{\ell}$. Consequently, a no-false-positive model cannot recognize a connected pair with distance $3^{L}$.
In both cases one can construct graphs with diameters $3^{L}$ on which $\mathsf{TF}_{\Theta}^{L}$ is not perfect. ∎
@@ 4.3 Training Dynamics
[…]
Figure 2: Capacity of Disentangled Transformers. We train 2-layer (left) and 3-layer (right) disentangled transformers on $\mathsf{ER}(n=24)$ and $\mathsf{ER}(n=64)$ graphs respectively. When evaluated on hold-out sets, both models can only make reliable predictions ($\geq 99\%$ accuracy) on node pairs $u,v$ if and only if $d_{G}(u,v)\leq 3^{L}$. It resonates with our theoretical observations in Theorem 4.4.
If a capable 2-layer Transformer is able to perfectly predict connectivity up to path length $3^{2}=9$, and the $\mathsf{2Chain}(n=20,k=10)$ dataset does not contain longer paths, why didn’t the Transformer model in §3.3 learn the algorithm? In this section, we show that this is because the training distribution contains too many graphs beyond the $3^{L}$ capacity, and those samples reward a global shortcut over algorithm. Equipped with Theorem 4.6, we can analyze the gradient dynamics in the two-channel parameterization (a superposition of heuristic and algorithmic channels). Theorems C.5 and C.9 give a simple criterion made possible by the exact $3^{L}$ characterization: if within-capacity pairs dominate, the algorithmic channel wins; if beyond-capacity pairs prevail, the shortcut wins.
Parameterizing Model Weights. We first need to formalize the criteria that define a “good” $\mathsf{TF}_{\Theta}^{L}$, especially if $\mathsf{ER}(n,p)$ exceeds the model’s capacity described in Theorem 4.4. The first notion is equivariance. Our data and targets are symmetric under node relabeling, in the sense that the ground truth mapping that maps adjacency matrices $A$ to connectivity matrices $R$ also maps $PAP^{\top}\mapsto PRP^{\top}$ for any permutation matrix $P$. It is therefore natural to demand that a “good” model’s output transform in the same way, i.e., $P\mathsf{TF}_{\Theta}^{L}(A)P^{\top}=\mathsf{TF}_{\Theta}^{L}(PAP^{\top})$. We use a generalized notion described in Appendix C.1. In addition, models attaining the tight capacity bound do so by learning the powering algorithm, so a reasonable second constraint is to enforce this: we require that $\mathsf{TF}_{\Theta}^{L}(A)$ and $A^{3^{L}}$ be identically supported, i.e., $\mathrm{supp}(\mathsf{TF}_{\Theta}^{L}(A))=\mathrm{supp}(A^{3^{L}})$. In Theorem C.2 we pin down the exact characterization of $W_{\ell}$ under these two conditions: the $(2^{\ell}n)\times(2^{\ell}n)$ matrix $W_{\ell}+W_{\ell}^{\top}$ must equal the Kronecker product of some $(2^{\ell}\times 2^{\ell})$ matrix $\Lambda_{\ell}$ with $I_{n}$.
For our analyses, however, we use a stronger assumption of layerwise attention equivariance, defined and characterized as follows:
@@
Suppose an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ has non-negative weights, i.e., $W_{\ell}\geq 0$ for all $\ell$. Then $\mathsf{TF}_{\Theta}^{L}$ is layer-wise permutation equivariant, i.e., for each $\ell$ and any hidden states $h\in\mathbb{R}^{n\times d_{\ell-1}}$,
@@
if and only if $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$ for some $A_{\ell},B_{\ell}\in\mathbb{R}^{2^{\ell}\times 2^{\ell}}$ for all $\ell$, where $\otimes$ denotes the Kronecker product and $J_{n}=\mathbf{1}\mathbf{1}^{T}$ the all-ones $(n\times n)$ matrix.
Sketch of proof.
Sufficiency is immediate: If $W_{\ell}$ admits this form, then conjugating by any node permutation $P$ leaves both factors invariant, as $PI_{n}P^{\top}=I_{n}$ and $PJ_{n}P^{\top}=J_{n}$.
For necessity, the key observation is that with ReLU inactive due to non-negativity assumption, the layer map becomes bilinear in $h$. With algebra, the equivariance assumption can be shown to imply a conjugation-alike identity on weights: Writing $\sigma(P)=I_{K_{\ell-1}}\otimes P$ and $\Delta=\sigma(P)W_{\ell}\sigma(P)^{\top}-W_{\ell}$, the following must hold:
$h\Delta h^{\top}h\sigma(P)=0\qquad\text{ for all }h\geq 0.$
We then argue that this forces $\Delta=0$, i.e., $W_{\ell}$ is conjugation-invariant, by testing the above equation with special matrices $P$. Next, we inspect small blocks $W_{\ell}[u,v]$ of size $n\times n$ in $W_{\ell}\in\mathbb{R}^{(2^{\ell}n)\times(2^{\ell}n)}$, and argue that $W_{\ell}[u,v]$ must commute with all permutations $P$. This forces each $W_{\ell}[u,v]$ to lie in $\mathrm{span}(I_{n},J_{n})$. Aggregating all subblocks, $W_{\ell}$ can therefore be decomposed as $A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$. ∎
This form is closed under gradients (Theorem C.4). As a result, the training dynamics yield a clean characterization via two channels: the algorithmic, powering mechanism via $A_{\ell}$, and the degree-counting heuristic via $B_{\ell}$, detailed immediately below. Importantly, we validate the layerwise attention equivariance assumption by showing models trained on these graphs rapidly display (approximate) layerwise equivariance in their hidden states (Figure 8). In addition, we show applicability of our training dynamics characterization on models not forced to be equivariant in §5.
@@
The Two-Channel Regime. Under the conditions of Theorem 4.6, each layer weight splits as $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$. This yields two functionally distinct channels (cf. equation 11):
  • The algorithmic $I_{n}$-channel ($A_{\ell}\otimes I_{n}$). It implements matrix powering algorithm: across layers one realizes combinations of $A^{k}$ (“$k$-hop” features), and the final readout layer combines them to compute $\sum_{j=1}^{\operatorname{diam}(G)}\alpha_{j}A^{j}$ in Theorem 4.3 and achieves the capacity bound in Theorem 4.4.
  • The heuristic $J_{n}$-channel ($B_{\ell}\otimes J_{n}$). The factor $J_{n}=\mathbf{1}\mathbf{1}^{\top}$ is rank-$1$ and satisfies, for vector $x$,
    $J_{n}x=(\mathbf{1}^{\top}x)\,\mathbf{1},\qquad AJ_{n}=(A\mathbf{1})\mathbf{1}^{\top}=\mathbf{d}\,\mathbf{1}^{\top},\qquad J_{n}A=\mathbf{1}\,(\mathbf{1}^{\top}A)=\mathbf{1}\,\mathbf{d}^{\top},$
    and more generally $A^{k}J_{n}=(A^{k}\mathbf{1})\mathbf{1}^{\top}$. Thus any use of the $J_{n}$-channel reduces to functions of $\mathbf{d}^{(k)}\ :=\ A^{k}\mathbf{1}$, the $k$-walk degree vectors (with $\mathbf{d}^{(1)}=\mathbf{d}$ the usual degree vector with $\mathbf{d}_{i}=\sum_{j=1}^{n}A_{i,j}$ the degree of node $i$.). Consequently, $B_{\ell}\otimes J_{n}$ contributes degree-counting heuristics and their $k$-walk generalizations, shared across nodes via broadcast.
Equipped with this algorithm-heuristic dual-channel view, we can analyze the training dynamics by tracking the two parameters $A_{\ell}$ and $B_{\ell}$. Before that, we make certain mild assumptions on training data distribution, model parameterization and loss objectives.
@@
When analyzing training dynamics of the disentangled transformers, we make the following assumptions
@@
    Data Distribution. Let $\mathsf{ER}(n,p)$ be the Erdős-Rényi distribution with edge-probability $p\in(0,1)$. Assume $\mathbb{P}_{G\sim\mathsf{ER}(n,p)}\{G\ \text{is disconnected}\}$ is bounded away from 0.
@@
    Nonnegativity & Equivariant Parameterization. For each layer $\ell$, we assume $W_{\ell}\geq 0$ and analyze the permutation-equivariant family $\mathrm{span}\{I_{n},J_{n}\}$ on the node side, i.e.,
    $W_{\ell}\;=\;A_{\ell}\otimes I_{n}\;+\;B_{\ell}\otimes J_{n},\qquad A_{\ell},B_{\ell}\in\mathbb{R}^{2^{\ell}\times 2^{\ell}}.$ (3)
@@
    Surrogate Loss. Given scores $Z:=\mathsf{TF}_{\Theta}^{L}(\cdot)\in\mathbb{R}_{\geq 0}^{n\times n}$, define the link $\phi(z):=1-e^{-\alpha z}$ with $\alpha>0$$\star$$\star\starEqual$ Contributionthe entrywise Bernoulli cross-entropy with respect to the connectivity matrix $R$ is
    $\mathcal{L}(Z;R)\ :=\ -\sum_{i,j}\big(R_{i,j}\log\phi(Z_{i,j})+(1-R_{i,j})\log(1-\phi(Z_{i,j}))\big).$ (4)
    Its gradient with respect to $Z$ is $\frac{\partial\mathcal{L}}{\partial Z}\ =\ \alpha\,(1-R/\phi(Z))\ \in\ \mathbb{R}^{n\times n}.$
[…]
Temporal Evolution. Roughly speaking, training consists of two phases.
@@
Phase 1: Both channels pick up easy examples. In early updates, both channels quickly ramp up mass because there are plenty of within-component, within-capacity pairs. Concretely, the local $I$-channel composes neighborhood information, while the global $J$-channel can also boost under-predicted positives without facing much penalty (Section C.3). Phase 1 ends once those easy connected pairs are mostly saturated, at which point false positives on cross-component terms become decisive, and the data-driven dichotomy of Phase 2 takes over. Overall, Phase 1 is transient: in Figure 3 (left), it only occupies around $2\cdot 10^{2}$ steps out of $10^{4}$ total.
Phase 2: Data determines which channel wins. Once in this regime, the growth of $B_{\ell}$ is determined by the population-level balance (see Theorem C.5 for full details and notations): informally,
$\text{Derivative of }B_{\ell}=\alpha\cdot\mathbb{E}\Bigg[\underbrace{\sum_{R_{i,j}=0}D_{i,j}}_{\begin{subarray}{c}\text{penalty from}\\ \text{cross component}\end{subarray}}-\underbrace{\sum_{R_{i,j}=1}\frac{1-\phi(Z_{i,j})}{\phi(Z_{i,j})}D_{i,j}}_{\begin{subarray}{c}\text{weighted reward on}\\ \text{under-predicted positives}\end{subarray}}\Bigg].$ (5)
There are two outcomes, depending on the sign of equation 5. If batches carry a significant mass of disconnected, within-capacity graphs, cross-component penalty dominates and suppresses the $J$-channel. Under these conditions, the only KKT stationary point in the $B$-coordinates is then $B_{\ell}=0$ (Theorem C.5(ii)), effectively only implementing the powering algorithm, making the model largely algorithmic. For instance, on Figure 4 (right, restricted), this shows up over the middle-to-late training window as the $A$-share climbs toward $1$ while the $B$-share decays to 0. Conversely, if the batches are rich in connected, beyond-capacity graphs, training instead promotes the $J$-channel, activating the dense degree-style, making the model largely heuristic.
A helpful way to interpret Phase 2 is via a mixture of graph types. The per-sample results (Theorem C.9) say: within-capacity graphs reward the local $I$-channel and, when disconnected, penalize any (falsely) activated $J$-channel; large-diameter connected graphs do the opposite and push $J$-channel up. Aggregating over the data distribution leads to the conclusion that the outcome of Phase 2 depends on the fraction of beyond-capacity connected graphs (Section C.4). Figure 6 visualizes the effects of various mixtures.
[…]
[width=]figures/training_dynamics/training_dynamics.pdf
Figure 3: Training Dynamics of Disentangled Transformers. We train an 1-layer disentangled transformer on graphs from $\mathsf{ER}(n=8,p=0.2)$ distribution. Weight $W$ will approximately approach to $A\otimes I_{n}+B\otimes J_{n}$ form. (Left) There are two major phases during training, where during Phase 1, model focuses on learning the equivariant parameterizations so both $I$ and $J$ channel’s share of energy in $W$ increases, and during Phase 2, the algorithmic $I$-channel is promoted and the heuristic $J$-channel is suppressed. (Right) Visualization of the learned weights during training and its projection to the closest $\hat{W}=\hat{A}\otimes I_{n}+\hat{B}\otimes J_{n}$ form.
[…]
[width=]figures/restrict_diam/restrict_diam_1layer_distf.pdf
Figure 4: Following insights from Theorems C.5 and C.9, we repeat the same experiment setup as in Figure 3 but only training on within-capacity graphs (see Section 4.2 and the left figure). As shown in the solid lines in the right figure, restricting training samples by capacity pushes the energy share of the algorithmic mechanism (the $A\otimes I_{n}$ channel) further to nearly 100% in the weight $W$. It simultaneously prevents the growth of the heuristic portion (the $B\otimes J_{n}$ channel).

### v2

@@ 4.2 Expressivity and Capacity
If a 2-layer Transformer fails to generalize in §3.3, should we attribute this to the architecture’s expressivity? We argue not. Theorem 4.3 shows that an $L$-layer Disentangled Transformer can implement the matrix powering algorithm and is perfect on graphs of diameter at most $3^{L}$. Moreover, Theorem 4.5 shows this $3^{L}$ threshold is tight and exact [^1] Because the $\mathsf{2Chain}(n=20,k=10)$ graphs have maximum path length $9$, they should be theoretically learnable by $2$-layer models, unlike in §3.3.. To make this precise, we first formalize graph distance and diameter in Definition 4.2.
@@
Throughout, we define the diameter of a graph, denoted $\mathrm{diam}(G)$, to be the maximum diameter among its connected components. Note this differs from the common convention on disconnected graphs, where the latter sets $\max_{u,v}d_{G}(u,v)=\infty$; ours is always finite.
@@
Theorem 4.3 (Expressivity of $\mathsf{TF}_{\Theta}^{L}$).
@@
For all $\ell$, setting $W_{\ell}=I_{d_{\ell-1}}$ suffices. These choices of weights implement the matrix powering algorithm $\sum_{j=0}^{\operatorname{diam}(G)}\alpha_{j}A^{j}$ with positive coefficients $\alpha_{j}$. ∎
We next show a capacity bound that reveals the model’s inherent limitations. We prove a tight, non-asymptotic upper bound on the graph diameter an $L$-layer model can handle, linking model depth directly to instance difficulty. To state this precisely, we define capacity as follows.
Definition 4.4 (Model Capacity).
The capacity of an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ is the largest integer $d$ such that there exist weights achieving perfect predictions on every graph $G$ with $\operatorname{diam}(G)\leq d$.
Informally speaking, the capacity of an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ with nonnegative weights is $3^{L}$. We formalize this claim as follows.
Theorem 4.5 (Capacity of $\mathsf{TF}_{\Theta}^{L}$).
Let $\mathsf{TF}_{\Theta}^{L}$ be an $L$-layer Disentangled Transformer on $n=\Omega(3^{L})$ nodes [^2] In particular, taking $n\geq(7/3)\cdot 3^{L}+2$ suffices.. Further assume that the weights $W_{\ell}\geq 0$ for each $\ell$. Then there exists a graph $G$ on which $\mathsf{TF}_{\Theta}^{L}(A)$ does not equal the connectivity matrix $R$. Further, $G$ has diameter $3^{L}+1$. In other words, $\mathsf{TF}_{\Theta}^{L}$ cannot master graph connectivity beyond path length $3^{L}$, and they cannot fully solve the task on $\Omega(3^{L})$-node graphs.
Figure 2: Capacity of Disentangled Transformers. We train 2-layer (top) and 3-layer (bottom) Disentangled Transformers on $\mathsf{ER}(n=24)$ and $\mathsf{ER}(n=64)$ graphs respectively. When evaluated on hold-out sets, both models can only make reliable predictions ($\geq 99\%$ accuracy) on node pairs $u,v$ if and only if $d_{G}(u,v)\leq 3^{L}$. These findings resonate with our theoretical observations in Theorem 4.5.
@@
Case 1 (False positive occurs Lemma B.1). Suppose for a graph $H$, an intermediate layer contains a false positive in its hidden states. That is, for two disconnected nodes $u,v$, the corresponding $(u,v)$ entry in some $\ell^{\text{th}}$ layer hidden states is positive (and this will propagate to the final output, whereas the $(u,v)$ entry in the ground-truth connectivity entry isn’t). We isolate two sets of nodes that contribute to this false positive by backtracking the computation DAG. By making appropriate changes, we argue for the existence of a graph $G$ that (i) preserves this false positive on $(u,v)$ and (ii) contains a manually created path of length $3^{L}+1$.
Case 2 (No false positives Lemma B.2). Suppose now that no intermediate layer has false positives for any $n$-node graph. We show that “information” spreads no faster than $3^{L}$ so that it never predicts “Yes” on node pairs with distance beyond $3^{L}$. We first apply the no-false-positives assumption on the empty (self-loops-only) graph. Inductively, each column of each hidden states is supported on exactly one row, which ranges from $1$ to $n$. This naturally gives a “label” for each column in each hidden states. The crux of the proof is to inductively show that at layer $\ell$, two columns can “share” information, thereby creating a positive score on $(u,v)$, only if their labels, interpreted as graph nodes, are within distance $3^{\ell}$. Consequently, a no-false-positive model cannot recognize a connected pair with distance $3^{L}+1$.
In both cases one can construct graphs with diameters $3^{L}+1$ on which $\mathsf{TF}_{\Theta}^{L}$ is not perfect. ∎
@@
Overloading the notation, we say $G$ is a within-capacity graph if $\mathrm{diam}(G)\leq 3^{L}$ and beyond-capacity otherwise.
@@ 4.3 Training Dynamics
Figure 3: Training Dynamics of Disentangled Transformers. We train a 1-layer Disentangled Transformer on graphs from $\mathsf{ER}(n=8,p=0.2)$ distribution. Weight $W$ will approximately approach to $A\otimes I_{n}+B\otimes J_{n}$ form. (Top) There are two major phases during training, where during Phase 1, model focuses on learning the equivariant parameterizations so both $I$ and $J$ channel’s share of energy (see §5.1) in $W$ increases, and during Phase 2, the algorithmic $I$-channel is promoted and the heuristic $J$-channel is suppressed. (Bottom) Visualization of the learned weights and its projection to the closest $\hat{W}=\hat{A}\otimes I_{n}+\hat{B}\otimes J_{n}$ form.
If a capable 2-layer Transformer is able to perfectly predict connectivity up to path length $3^{2}=9$, and the $\mathsf{2Chain}(n=20,k=10)$ dataset does not contain longer paths, why didn’t the Transformer model in §3.3 learn the algorithm? In this section, we show that this is because the training distribution contains too many graphs beyond the $3^{L}$ capacity, and those samples reward a learning a global heuristic over an algorithm. Equipped with Theorem 4.7, we can analyze the gradient dynamics in the two-channel parameterization (a superposition of heuristic and algorithmic channels). Theorem C.5 and Theorem C.9 give a simple criterion made possible by the exact $3^{L}$ characterization: if within-capacity pairs dominate, the algorithmic channel wins; if beyond-capacity pairs prevail, the heuristic wins.
Parameterizing Model Weights. To analyze the gradient dynamics, we first identify the relevant parameter space. Our data and targets are symmetric under node relabeling: the ground truth mapping maps $A\mapsto R$ and $PAP^{\top}\mapsto PRP^{\top}$ for any permutation $P$. We also empirically observe that models trained from scratch on these graphs rapidly converge to a layerwise equivariant state (Figure 7). Based on these observations, we analyze the dynamics within the subspace of layerwise permutation-equivariant weights. The following theorem formally defines layerwise equivariance and characterizes exactly what weights look like.
@@
Suppose an $L$-layer Disentangled Transformer $\mathsf{TF}_{\Theta}^{L}$ has non-negative weights. Let $K_{\ell-1}=2^{\ell}$. Then $\mathsf{TF}_{\Theta}^{L}$ is layer-wise permutation equivariant, i.e., for each $\ell$, any hidden states $h\in\mathbb{R}^{n\times d_{\ell-1}}$, and any permutation $P\in S_{n}$,
@@
if and only if each layer weight $W_{\ell}$ decomposes as $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$ for some $A_{\ell},B_{\ell}\in\mathbb{R}^{2^{\ell}\times 2^{\ell}}$ for all $\ell$, where $\otimes$ denotes the Kronecker product and $J_{n}=\mathbf{1}\mathbf{1}^{T}$ the all-ones $(n\times n)$ matrix.
It immediately follows that this subspace contains a canonical algorithmic solution (e.g. the identity construction $W=I$ used in Theorem 4.3). Furthermore, Theorem C.2 (in Appendix) shows that for any capacity-reaching model, the symmetric component of the weights, which drives the attention mechanism, must lie purely in the $I_{n}$-channel. Finally, algebraically this parameterization is closed under gradients (Theorem C.4), and this allows us to decompose the learning process into the competition between two functionally distinct channels, discussed next.
@@
Under the conditions of Theorem 4.7, each layer weight decomposes as $W_{\ell}=A_{\ell}\otimes I_{n}+B_{\ell}\otimes J_{n}$. This decomposition separates the model’s computation into two functionally distinct channels with different roles.
The $I_{n}$-channel ($A_{\ell}\otimes I_{n}$) compute algorithms.
The term $A_{\ell}\otimes I_{n}$ preserves locality: when applied within the attention mechanism, it only combines features between nodes that share graph neighbors. Across layers, this channel composes multi-hop information by effectively computing powers of the adjacency matrix $A$. After $L$ layers, the readout aggregates a weighted sum $\sum_{j=0}^{3^{L}}\alpha_{j}A^{j}$ with nonnegative coefficients, which is precisely the matrix powering algorithm that solves connectivity (Theorem 4.3).
The $J_{n}$-channel ($B_{\ell}\otimes J_{n}$) collect heuristics.
The term $B_{\ell}\otimes J_{n}$ broadcasts information globally, ignoring graph structure. To see why, observe that $J_{n}$ is rank-one: for any vector $x\in\mathbb{R}^{n}$, we have $J_{n}x=(\mathbf{1}^{\top}x)\mathbf{1}$, which computes the sum of entries and broadcasts it uniformly to all nodes. When composed with the adjacency matrix, $AJ_{n}=\mathbf{d}\mathbf{1}^{\top}$ where $\mathbf{d}=A\mathbf{1}$ is the degree vector. Thus, this channel computes global statistics, specifically products of node degrees and their higher-order generalizations. Such statistics correlate with connectivity on dense random graphs but fail on adversarial instances. For example, two disjoint cliques both have high-degree nodes, yet no edge connects them.
Training Dynamics. Under this algorithmic-herustic dual-channel view, we can track the evolution of the two parameters, $A_{\ell}$ and $B_{\ell}$. To rigorously analyze the gradient descents, we adopt the following assumptions.
@@
We make assumptions on data distribution, model parameterization and the training loss.
@@
    Data Distribution. Let $\mathsf{ER}(n,p)$ be the Erdős-Rényi distribution with edge-probability $p\in(0,1)$. Then $\mathbb{P}_{G\sim\mathsf{ER}(n,p)}\{G\ \text{is disconnected}\}>0$.
@@
    Nonnegativity & Equivariant Parameterization. For each layer $\ell$, assume $W_{\ell}\geq 0$ can be decomposed as $W_{\ell}\;=\;A_{\ell}\otimes I_{n}\;+\;B_{\ell}\otimes J_{n}$ for some $A_{\ell},B_{\ell}\in\mathbb{R}^{2^{\ell}\times 2^{\ell}}.$
@@
    Surrogate Loss. Given scores $Z:=\mathsf{TF}_{\Theta}^{L}(\cdot)\in\mathbb{R}_{\geq 0}^{n\times n}$, define the link $\phi(z):=1-e^{-\alpha z}$ with $\alpha>0$; [^3] It is possible that $R_{i,j}=1$ while $Z_{i,j}=0$, resulting in undefined gradient $\partial\mathcal{L}/\partial Z$. To circumvent this, we approximate via $\phi_{\epsilon}=1-(1-\epsilon)e^{-\alpha z}$. All subsequent analyses hold verbatim by replacing $\phi$ with $\phi_{\epsilon}$.the entrywise cross-entropy with respect to the connectivity matrix $R$ is $\mathcal{L}(Z;R)\ :=\ -\sum_{i,j}\big(R_{i,j}\log\phi(Z_{i,j})+(1-R_{i,j})\log(1-\phi(Z_{i,j}))\big).$ Its gradient with respect to $Z$ is $\frac{\partial\mathcal{L}}{\partial Z}\ =\ \alpha\,(1-R/\phi(Z))\ \in\ \mathbb{R}^{n\times n}.$
Under these assumptions, we can characterize the convergence and limiting behavior of gradient descent.
Theorem 4.9 (Convergence to KKT Points).
Let $\mathcal{R}(\Theta):=\mathbb{E}_{G\sim\mathsf{ER}(n,p)}[\mathcal{L}(\mathsf{TF}_{\Theta}^{L}(A_{G});R_{G})]$ denote the population risk. For $\lambda>0$, define the regularized objective $\mathcal{R}_{\lambda}(\Theta):=\mathcal{R}(\Theta)+\frac{\lambda}{2}\|\Theta\|_{F}^{2}$. Let $\mathcal{C}:=\{(A_{\ell},B_{\ell})_{\ell}:A_{\ell}\geq 0,B_{\ell}\geq 0,\forall\ell\}$ denote the constraint set, and consider the sequence $\{\Theta^{(k)}\}_{k\geq 0}$ generated by projected gradient descent on $\mathcal{R}_{\lambda}$:
$\Theta^{(k+1)}=\Pi_{\mathcal{C}}\left(\Theta^{(k)}-\eta\nabla\mathcal{R}_{\lambda}(\Theta^{(k)})\right),$ (1)
with step size $\eta>0$ sufficiently small and initialization $\Theta^{(0)}\in\mathcal{C}$ of the form $W_{\ell}=A_{\ell}\otimes I+B_{\ell}\otimes J$. Then every limit point $\Theta^{*}_{\lambda}\in\mathcal{C}$ satisfies the KKT conditions:
$\nabla_{B_{\ell}}\mathcal{R}(\Theta^{*}_{\lambda})+\lambda B_{\ell}^{*}\geq 0,\quad B_{\ell}^{*}\geq 0,$ (2)
$\left(\nabla_{B_{\ell}}\mathcal{R}(\Theta^{*}_{\lambda})+\lambda B_{\ell}^{*}\right)\odot B_{\ell}^{*}=0,$
and analogously for $A_{\ell}^{*}$. Moreover, the iterates converge to a KKT point at the standard $\mathcal{O}(1/\epsilon)$ rate for projected gradient descent.
The proof adapts standard convergence analysis for projected gradient descent on smooth nonconvex functions (Bertsekas, 1997; Beck, 2017). See Section C.5 for details. In the next result, we analyze the structure of the KKT points of the objective.
Theorem 4.10 (Learning the Algorithm, informal version of Theorem C.5 and Corollary C.20).
Under suitable conditions where the gradient penalty from disconnected graphs outweighs the gradient reward from connected graphs, the only KKT-compliant value for the heuristic channel is $B_{\ell}^{*}=0$, i.e., the model converges to learning the fully algorithmic matrix-powering algorithm.
While Theorem 4.9 guarantees that training converges to limit points $\Theta_{\lambda}^{*}$ of the objective $\mathcal{R}_{\lambda}(\Theta)$, Theorem 4.10 further characterizes a sufficient condition for algorithmic alignment. The formal versions for the un-regularized and regularized objectives can be found at Theorem C.5 and Corollary C.20, respectively. With convergence guaranteed, our analysis of gradient descents reveals that the training process consists of two distinct phases:
@@
Phase 1: Both channels pick up easy examples. In early updates, both channels quickly ramp up mass because there are plenty of within-component, within-capacity pairs. Concretely, the local $I$-channel composes neighborhood information, while the global $J$-channel can also boost under-predicted positives without facing much penalty (Remark C.7). Phase $1$ is transient and ends once those easily connected pairs are mostly saturated. In Figure 3 (top), it only occupies around $2\cdot 10^{2}$ steps out of $10^{4}$ total.
Phase 2: Data determines which channel wins. Once in this regime, the growth of $B_{\ell}$ is determined by the population-level balance (see Theorem C.5 for full definitions and details). Informally speaking, the derivative of $B_{\ell}$ is driven by a competition between a suppression force and a promotion force. The suppression force arises from disconnected graphs, where the heuristic generates false positives, and they push $B_{\ell}$ to zero. The promotion force arises from connected graphs, where the heuristic helps minimize loss by correctly identifying connected pairs, generating a reward gradient.
There are two outcomes, best understood through the distinction between within-capacity and beyond-capacity graphs (Theorem C.9). If batches are dominated by within-capacity graphs ($\operatorname{diam}(G)\leq 3^{L}$), the algorithmic $I$-channel is consistently rewarded. Disconnected graphs in this regime penalize the heuristic $J$-channel for predicting false connections, driving $B_{\ell}\to 0$ (Theorem C.5(ii)) and leaving only the algorithmic solution. Conversely, if the distribution contains a significant share of connected beyond-capacity graphs ($\operatorname{diam}(G)>3^{L}$), the algorithmic channel cannot bridge the distance. These samples force the model to rely on the global $J$-channel to minimize loss, promoting the degree-counting shortcut. Thus, the final learned solution depends directly on the proportion of beyond-capacity connected graphs in the training distribution (Remark C.13), as visualized in Figure 5.